Use the properties of logarithms to solve the equation for x.
log2(5x − 27) = 3

Answer by lwsshak3(11628)

solve for x
log2(5x − 27) = 3
convert to exponential form: base(2) raised to log of number(3)=number(5x-27)
2^3=5x-27=8
5x=8+27=35
x=7